# She’s Got Her Mother’s Hair: End-to-End Collapsar Simulations Unveil the Origin of Black Holes’ Magnetic Field

### v1

@@ 5 Discussion
Alternatively, the strong magnetic field threading the BH could be advected during the free-fall collapse while maintaining flux-freezing. We demonstrate that under the TSD mechanism, stellar evolution models suggest the magnetic field loops are too small to add up coherently. The resulting flux on the BH falls short by orders of magnitude compared to that needed to power a GRB jet. The inability of both scenarios to generate a strong large-scale poloidal magnetic field poses significant challenges in understanding the origin of the BH’s magnetic field.
Our stellar evolution models reveal that rapidly rotating progenitors associated with GRBs form a PNS with an accretion disk prior to BH formation. During the PNS phase, the convective dynamo within the PNS amplifies its internal magnetic field to $B\approx 10^{15}\,\,{{\rm G}}$. As the PNS reaches a critical mass and collapses into a BH, some of its magnetic field is inherited by the BH, depending on the relative length of the balding timescale and the viscous time. We find that, for a characteristic mass accretion rate and PNS field, the viscous time is shorter than the balding timescale, allowing the field lines to rearrange and enabling the nascent BH to retain the flux. Conversely, if the mass accretion rate onto the PNS is very low and the magnetic field is extremely strong, the magnetic footprints lie outside the PNS, resulting in most of the flux being lost to reconnection during the collapse. Nonetheless, even if the balding time is too short, causing the bulk of the PNS flux to be lost during the collapse, the BH is likely to inherit the highly magnetized PNS accretion disk. This disk can serve as a seed for further amplification of the magnetic field, ultimately leading to the launch of a BZ jet. This suggests that even if the BH flux is generated by an accretion disk, it is likely that the disk is an inherited, highly magnetized PNS disk. Namely, there is no necessity for a newly formed accretion disk to gradually amplify the weaker fossil star fields through dynamo processes.
@@
To determine whether the BH can sustain the launch of jets without an additional supply of magnetic flux, we conducted 3D GRMHD simulations of BHs with the expected magnetic flux from the PNS, embedded inside a star based on our stellar evolution models. Our findings indicate that an accretion disk is essential for the BH to retain its magnetic field, enabling a steady launch of relativistic jets throughout the simulations ($t>6\,\,{{\rm s}}$) with a roughly constant luminosity. Our jets are not in a MAD state, which requires higher magnetic flux and a lower mass accretion rate than current simulations predict. If no additional magnetic flux is accreted onto the BH, the BH will remain in a sub-MAD state. However, over time, the mass accretion rate is expected to decrease, increasing the dimensionless magnetic flux on the BH until it goes MAD. In scenarios where the disk is absent, the magnetic flux reconnects along the equator, preventing the BH from launching jets. Future numerical relativity simulations that follow the NS collapse into the BH will be conducted to provide a rigorous, self-consistent test of this scenario.
@@ 6 Big Picture Insights on gamma-ray bursts and associated supernovae
In this letter, we have explored rapidly rotating stars that support the necessary conditions for accretion disk formation, which is essential for GRB jets, and subsequently lead to the birth of rapidly spinning magnetars. Such proto-magnetars typically collapse into BHs within $\sim 1\,\,{{\rm s}}$. During this time, they release $\sim 10^{51}-10^{52}\,\,{{\rm erg}}$ into the collapsing star, potentially accounting for the excess energy observed in SNe Ic-BL associated with collapsar GRBs. However, in some cases, the magnetar winds may impede accretion sufficiently to prevent BH collapse. Consequently, the ms magnetar may remain active for an extended period, potentially powering Type I superluminous SNe (e.g., Kasen & Bildsten, 2010; Woosley, 2010; Nicholl et al., 2017; Vurm & Metzger, 2021).
@@
In contrast, more common stars may possess stronger, large-scale magnetic fields. These fields could be of fossil origin, or generated through dynamo action in early evolutionary phases (e.g. Donati & Landstreet, 2009; Fuller et al., 2015; Cantiello et al., 2016). The presence of such large-scale fields facilitates efficient angular momentum transport away from the core. As a result, these stars are expected to form slowly spinning magnetars, which might be associated with the SN engine.
@@
Following the collapse of the PNS into a BH, a BZ driven jet with a constant power of $P_{j}\sim 10^{49}\,\,{{\rm erg}}$ is launched into the collapsing star. Once this jet breaks out of the optically thick star, it can generate the prompt GRB emission. Both the constant jet power and its magnitude align with observations of typical GRBs (e.g., McBreen et al., 2002; Butler et al., 2010), where the GRB rapid variability naturally emerges from the jet-star interaction (Gottlieb et al., 2019). Nonetheless, GRBs exhibit a wide spectrum of luminosities (e.g., Wanderman & Piran, 2010). Within our model, the jet power varies with the BH spin and the magnetic flux. Namely, the range of GRB luminosities may stem from variations in mass accretion rates and magnetic fluxes on the PNS, which in turn determine the resulting natal BH spin and dipolar field. During the jet activity, the BH is expected to enter the MAD state, leading to accelerated BH spin-down. Therefore, the end-point of the jet launching is likely governed by either the decreasing mass accretion rate, the spin-down of the BH, or a combination of both factors [see Eq. (6)].
@@
Finally, the current paradigm posits two stringent requirements for the progenitor stars of GRBs: they must simultaneously maintain strong magnetic fields and rapid rotation (e.g., Symbalisty, 1984; Metzger et al., 2008; Gottlieb et al., 2022a). However, strong magnetic fields carry away angular momentum from the star through magnetized winds, making the co-existence of a strong field and rapid rotation challenging. If BHs inherit their magnetic field from PNSs, as proposed here, these requirements are alleviated, making rapid rotation of the star the primary requirement for jet launching. This suggests the following: (1) relativistic jets might form more readily and, hence, be more common. Furthermore, it also suggests that the majority of collapsar BH accretion disks will be accompanied by magnetized jets; (2) BHs formed through direct collapse cannot achieve strong magnetic fields and thus cannot produce jets, implying that a PNS phase is crucial for typical collapsar GRBs. Other GRB classes, such as ultra-long GRBs, may require a different launching mechanism, suggesting they are fundamentally distinct from ordinary GRBs.
@@
We are grateful to Lucy Reading-Ikkanda/Simons Foundation for designing Figure 1. We thank Amir Levinson, Wenbin Lu, Jonatan Jacquemin-Ide, Elias Most, Jim Fuller, and Stan Woosley for their helpful comments. OG and JAG are supported by the Flatiron Research Fellowship. BDM was supported in part by the National Science Foundation (grant No. AST-2009255) and by the NASA Fermi Guest Investigator Program (grant No. 80NSSC22K1574). The Flatiron Institute is supported by the Simons Foundation.
@@ Appendix A Evolution of the stellar progenitor
The minimum radius $\sim{}0.5\,R_{\odot}$ is reached roughly at the onset of Neon core ignition and the evolution afterward is particularly sensitive to the treatment of convection, numerical resolution, and nuclear physics. To improve the numerical convergence, we increase the resolution in Lagrangian mass coordinate based on the radiative, adiabatic and mean molecular weight gradients across the local pressure scale height with a custom other_mesh_fcn_data (see run_star_extras.f90 available at doi.org/10.5281/zenodo.12193630).
@@
The left panels of Fig. 6 show a resolution test on the structure at the onset of core-collapse. We recomputed the model from oxygen core depletion (defined as the first time during the evolution after the end of hydrogen core burning when the central mass fraction of ${}^{16}\mathrm{O}$ is less than 0.1 and the mass fractions of ${}^{12}\mathrm{C}$ and ${}^{4}\mathrm{He}$ are below 0.001 and 0.005, respectively) with increased spatial and temporal resolution (mesh_delta_coeff, mesh_delta_coeff_for_highT, and mesh_time_coeff set to 0.75, cf. 1.0 in our standard setup). This produces the dashed orange lines in Fig. 6. The legend in the top right panel gives the number of mesh points in the profiles at the onset of core collapse. The blue model corresponds to the one discussed in the main text.
Figure 6: Resolution test for $40\,M_{\odot}$, $Z=0.001$ initially rotating at $0.6\omega_{\mathrm{crit}}$ model, which experiences rotationally induced chemically homogeneous evolution. The left panels show, from top to bottom, the specific angular momentum, density, temperature, and velocity profile as a function of radius at the onset of core-collapse. The right panel shows the Hertzsprung-Russell diagram of the Each dot marks $10^{5}$ years of evolution. The label indicates the number of mesh points at the onset of core collapse. The higher resolution model was recomputed post oxygen core depletion only.

### v2

@@ 5 Discussion
Alternatively, the strong magnetic field threading the BH could be advected during the free-fall collapse while maintaining flux freezing. We demonstrate that under the TSD mechanism, stellar evolution models suggest the magnetic field loops are too small to add up coherently. The resulting flux on the BH falls short by orders of magnitude compared to that needed to power a GRB jet. The inability of both scenarios to generate a strong large-scale poloidal magnetic field poses significant challenges in understanding the origin of the BH’s magnetic field.
Our stellar evolution models reveal that rapidly rotating progenitors associated with GRBs exhibit a large amount of angular momentum in the core, forming a PNS with an accretion disk prior to BH formation. During the PNS phase, the convective dynamo within the PNS amplifies its internal magnetic field to $B\approx 10^{15}\,\,{{\rm G}}$. As the PNS reaches a critical mass and collapses into a BH, some of its magnetic field is inherited by the BH, depending on the relative length of the balding timescale and the viscous time. We find that, for a characteristic mass accretion rate and PNS field, the viscous time is shorter than the balding timescale, allowing the field lines to rearrange and enabling the nascent BH to retain the flux. Conversely, if the mass accretion rate onto the PNS is very low and the magnetic field is extremely strong, the magnetic footprints lie outside the PNS, resulting in most of the flux being lost to reconnection during the collapse. Nonetheless, even if the balding time is too short, causing the bulk of the PNS flux to be lost during the collapse, the BH is likely to inherit the highly magnetized PNS accretion disk. This disk can serve as a seed for further amplification of the magnetic field, ultimately leading to the launch of a BZ jet. This suggests that even if the BH flux is generated by an accretion disk, it is likely that the disk is an inherited, highly magnetized PNS disk. Namely, there is no necessity for a newly formed accretion disk to gradually amplify the weaker fossil star fields through dynamo processes.
@@
To determine whether the BH can sustain the launch of jets without an additional supply of magnetic flux, we conducted 3D GRMHD simulations of BHs with the expected magnetic flux from the PNS embedded inside a star based on our stellar evolution models. Our findings indicate that an accretion disk is essential for the BH to retain its magnetic field, enabling a steady launch of relativistic jets throughout the simulations ($t>6\,\,{{\rm s}}$) with a roughly constant luminosity. Our jets are not in a MAD state, which requires higher magnetic flux and a lower mass accretion rate than current simulations predict. If no additional magnetic flux is accreted onto the BH, the BH will remain in a sub-MAD state. However, over time, the mass accretion rate is expected to decrease, increasing the dimensionless magnetic flux on the BH until it goes MAD. In scenarios where the disk is absent, the magnetic flux reconnects along the equator, preventing the BH from launching jets. Numerical relativity simulations that follow the PNS phase, the PNS-disk interaction, and the PNS collapse into a BH are crucial to provide a rigorous, self-consistent test of this scenario and will be conducted in a follow-up study.
@@ 6 Big Picture Insights on gamma-ray bursts and associated supernovae
In this letter, we have explored rapidly rotating stars that support the necessary conditions for accretion disk formation, which is essential for GRB jets, and subsequently lead to the birth of rapidly spinning magnetars. Such proto-magnetars typically collapse into BHs within $\sim 1\,\,{{\rm s}}$. During this time, they release $\sim 10^{51}-10^{52}\,\,{{\rm erg}}$ into the collapsing star, potentially accounting for the excess energy observed in Type Ic-BL SNe associated with collapsar GRBs. However, in some cases, the magnetar winds may impede accretion sufficiently to prevent BH collapse. Consequently, the ms magnetar may remain active for an extended period, potentially powering Type I superluminous SNe (e.g., Kasen & Bildsten, 2010; Woosley, 2010; Nicholl et al., 2017; Vurm & Metzger, 2021).
@@
In contrast, more common stars may possess stronger large-scale magnetic fields. These fields could be of fossil origin or generated through dynamo action in early evolutionary phases (e.g. Donati & Landstreet, 2009; Fuller et al., 2015; Cantiello et al., 2016). The presence of such large-scale fields facilitates efficient angular momentum transport away from the core. As a result, these stars are expected to form slowly spinning magnetars, which might be associated with the SN engine.
@@
Following the collapse of the PNS into a BH, a BZ driven jet with a constant power of $P_{j}\sim 10^{49}\,\,{{\rm erg}}$ is launched into the collapsing star. Once this jet breaks out of the optically thick star, it can generate the prompt GRB emission. Both the constant jet power and its magnitude align with observations of typical GRBs (e.g., McBreen et al., 2002; Butler et al., 2010), where the GRB rapid variability naturally emerges from the jet-star interaction (Gottlieb et al., 2019). Nonetheless, GRBs exhibit a wide spectrum of luminosities (e.g., Wanderman & Piran, 2010). Within our model, the jet power varies with the BH spin and the magnetic flux. Namely, the range of GRB luminosities may stem from variations in mass accretion rates and magnetic fluxes on the PNS, which in turn determine the resulting natal BH spin and dipolar field. During the jet activity, the BH is expected to enter the MAD state, leading to accelerated BH spin-down. Therefore, the end point of the jet launching is likely governed by either the decreasing mass accretion rate, the spin-down of the BH, or a combination of both factors [see Eq. (6)].
@@
Finally, the current paradigm posits two stringent requirements for the progenitor stars of GRBs: they must simultaneously maintain strong magnetic fields and rapid rotation (e.g., Symbalisty, 1984; Metzger et al., 2008; Gottlieb et al., 2022a). However, strong magnetic fields carry away angular momentum from the star through magnetized winds (e.g., Varma & Müller, 2023; Petitdemange et al., 2023a, b), making the coexistence of a strong field and rapid rotation challenging. If BHs inherit their magnetic field from PNSs, as proposed here, these requirements are alleviated, making rapid rotation of the star the primary requirement for jet launching. This suggests the following: (1) Relativistic jets might form more readily and, hence, be more common. Furthermore, it also suggests that the majority of collapsar BH accretion disks will be accompanied by magnetized jets; (2) BHs formed through direct collapse such as in pair-instability SNe cannot achieve strong magnetic fields and thus cannot produce jets, implying that a PNS phase is crucial for typical collapsar GRBs. Other GRB classes, such as ultra-long GRBs, may require a different launching mechanism, suggesting that they are fundamentally distinct from ordinary GRBs.
@@
We are grateful to Lucy Reading-Ikkanda/Simons Foundation for designing Figure 1. We thank Valentin Skoutnev, Amir Levinson, Wenbin Lu, Jonatan Jacquemin-Ide, Elias Most, Jim Fuller, Stan Woosley, and Adam Burrows for their helpful comments. MR thanks Aldana Grichener for useful feedback on the reproducibility package on zenodo. OG and JAG are supported by the Flatiron Research Fellowship. BDM was supported in part by the National Science Foundation (grant No. AST-2009255) and by the NASA Fermi Guest Investigator Program (grant No. 80NSSC22K1574). The Flatiron Institute is supported by the Simons Foundation. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC allocations m4603 (award NP-ERCAP0029085).
@@ Appendix A Evolution of the stellar progenitor
The minimum radius $\sim{}0.5\,R_{\odot}$ is reached roughly at the onset of Neon core ignition and the evolution afterward is particularly sensitive to the treatment of convection, numerical resolution, and nuclear physics. To improve the numerical convergence, we increase the resolution in Lagrangian mass coordinates based on the radiative, adiabatic and mean molecular weight gradients across the local pressure scale height with a custom other_mesh_fcn_data (see run_star_extras.f90 available at doi.org/10.5281/zenodo.12193630).
@@
The left panels of Fig. 6 show a resolution test on the structure at the onset of core-collapse. We recomputed the model from oxygen core depletion (defined as the first time during the evolution after the end of hydrogen core burning when the central mass fraction of ${}^{16}\mathrm{O}$ is less than 0.1 and the mass fractions of ${}^{12}\mathrm{C}$ and ${}^{4}\mathrm{He}$ are below 0.001 and 0.005, respectively) with increased spatial and temporal resolution (mesh_delta_coeff, mesh_delta_coeff_for_highT, and mesh_time_coeff set to 0.75, cf. 1.0 in our standard setup). This produces the dashed orange lines in Fig. 6. The legend in the right panel gives the number of mesh points in the profiles at the onset of core collapse. The blue model corresponds to the one discussed in the main text.
Figure 6: Resolution test for $40\,M_{\odot}$, $Z=0.001$ initially rotating at $0.6\omega_{\mathrm{crit}}$ model, which experiences rotationally induced chemically homogeneous evolution. The left panels show, from top to bottom, the specific angular momentum, density, temperature, and velocity profile as a function of radius at the onset of core-collapse. The right panel shows the Hertzsprung-Russell diagram, where each dot marks $10^{5}$ years of evolution. The label indicates the number of mesh points at the onset of core collapse. The higher resolution model was recomputed post oxygen core depletion only.